Find f_x(x,y), f xy(x,y), f x(1,3), and f y(-2,4) for the given function.


1
Expert's answer
2021-06-15T07:18:30-0400
"\ud835\udc67 = \ud835\udc53(\ud835\udc65, \ud835\udc66) = 3\ud835\udc65^3\ud835\udc66^2 \u2212 \ud835\udc65^2\ud835\udc66^3 + 4\ud835\udc65 + 9"


"f_x(x, y)=9x^2y^2-2xy^3+4"

"f_y(x, y)=6x^3y-3x^2y^2"




"f_{xx}(x, y)=18xy^2-2y^3"

"f_{xy}(x, y)=18x^2y-6xy^2=f_{yx}(x, y)"

"f_{yy}(x, y)=6x^3-6x^2y"



"f_x(1, 3)=9(1)^2(3)^2-2(1)(3)^3+4=31"

"f_y(-2, 4)=6(-2)^3(4)-3(-2)^2(4)^2=-384"